OpenStudy (ammarah):

Graph the equation. Identify the vertices , co-vertices, and foci of the ellipse.

OpenStudy (ammarah):



OpenStudy (ammarah):

@ganeshie8

OpenStudy (ammarah):

@whpalmer4

OpenStudy (ammarah):

@satellite73

OpenStudy (ammarah):

@nikato

OpenStudy (ammarah):

@agent0smith

ganeshie8 (ganeshie8):

its an ellipse

OpenStudy (ammarah):

wow that was hellpful but idk how to solve it like the whole a and b thing

ganeshie8 (ganeshie8):

since the number below x is less than the the number below y, u need to use last column in the chart

OpenStudy (ammarah):

really isnt it the fist clollumn

ganeshie8 (ganeshie8):

\(\large \frac{x^2}{16} + \frac{y^2}{36} = 1\) \(\large \frac{x^2}{4^2} + \frac{y^2}{6^2} = 1\)